OpenStudy (anonymous):

How do I create a perfect square trinomial? That can be factored multiple times. I need the process of creating my own in complete sentences.

OpenStudy (phi):

a perfect square is easy to create. pick a binomial such as (x+2) or (x-1) or (x-10) then multiply it by itself using FOIL (first, outer, inner, last) and combine like terms

OpenStudy (phi):

example (x+1)(x+1) = x*x +x*1 + 1*x + 1 = x^2 +2x +1

OpenStudy (anonymous):

a similar question was asked before...take a look http://openstudy.com/study#/updates/4eb94058e4b0e7092ba8415d it may help you!